OpenStudy (anonymous):

15. Choose the equation below that represents the line that passes through the point (7, –2) and has a slope of –3. (3 points)

OpenStudy (anonymous):

y - y1 = m(x-x1)

OpenStudy (anonymous):

m=3, y1 = -2, x1 = 7. Put it in the equation: y - y1=m(x-x1)

OpenStudy (anonymous):

sorry m = -3

OpenStudy (anonymous):

DO you need anymore help?

OpenStudy (anonymous):

Little bit

OpenStudy (anonymous):

The (y-v1)=m(x-x1) thing the people have been talking about is an equation of a line called point slope form where (x1,y1) is a point on the line and m is the slope of the line. Therefore with the given information, you can create an equation for the line using the point-slope form

OpenStudy (anonymous):

Try putting you point and slope into the equation and I will check

OpenStudy (anonymous):

Y- 2 = -3(x+7).

OpenStudy (anonymous):

Almost. Remember, it is (y-y1) and (x-x1) a minus sign. it would be (y+2)=-3(x-7)

OpenStudy (anonymous):

Oh thanks so much. Live saver.

OpenStudy (anonymous):

If you want to put it into a more familiar form, you can distribute to get y+2=-3x+21 y=-3x+19 and you're welcome

OpenStudy (anonymous):

See ya.